### v1

# Non-Hermitian Multipole Skin Effects Challenge Localization

Abstract
We study the effect of quenched disorder on the non-Hermitian skin effect in systems that conserve a U(1) charge and its associated multipole moments. In particular, we generalize the Hatano-Nelson argument for an Anderson transition in a disordered system with non-reciprocal hopping to the interacting case. When only U(1) charge is conserved, we show that there is a transition between a skin effect phase, in which charges are localized at a boundary, and a many-body localized phase, in which the charges are localized at random positions. In periodic boundary conditions, the skin effect gives way to a delocalized phase with a unidirectional current. If additional multipole moments like dipoles are conserved, we show that the non-Hermitian skin effect remains stable for any amount of disorder. As a result, under periodic boundary conditions, the system is always delocalized regardless of disorder strength.
Introduction.—The theory of Anderson localization is a milestone in condensed matter physics that describes how a disordered potential can trap non-interacting quantum particles Anderson (1958). This effect has been generalized to interacting systems—dubbed many-body localization (MBL)—and challenges our understanding of how the laws of statistical mechanics emerge in a closed system Basko et al. (2006); Gornyi et al. (2005); Nandkishore and Huse (2015); Abanin et al. (2019); Sierant et al. (2025). In practice, however, complete isolation is only approximate. Interactions between a quantum system and its environment can induce dissipation, decoherence, and wavefunction collapse, destroying unitarity. Many of the resulting effects can be captured by non-Hermitian Hamiltonians Xiujuan Zhang and Chen (2022); Yuto Ashida and Ueda (2020).
Non-Hermitian systems host a rich variety of behaviors that are not possible in their Hermitian counterparts, ranging from generalized topological phases to novel types of quantum criticality and entanglement transitions Rudner and Levitov (2009); Hu and Hughes (2011); Borgnia et al. (2020); Lee et al. (2019); Okuma et al. (2020); Gong et al. (2018); Yao and Wang (2018); Chen et al. (2020); Bergholtz et al. (2021); Tang et al. (2021); Li et al. (2018); *Li_Yao_19; Skinner et al. (2019); Zabalo et al. (2022); Chan et al. (2019); Turkeshi et al. (2021). One of the most celebrated phenomena is the non-Hermitian skin effect, where an extensive number of single-particle eigenstates are localized at the boundary of a lattice because of non-reciprocal hopping Hatano and Nelson (1996); *Hatano1997; *Hatano1998NonHermitian; Brouwer et al. (1997); Feinberg and Zee (1999). The skin effect emerges from non-Hermitian topology and realizes a contrasting bulk-boundary correspondence to that of Hermitian topological phases Lee et al. (2019); Okuma et al. (2020). Recently, the skin effect was generalized to systems that conserve multipole moments of a scalar U$(1)$ charge Gliozzi et al. (2024). Multipole conserving systems are intrinsically interacting, and the key signature of the multipole skin effect in an $m$-pole conserving system is eigenstates with an extensive $(m+1)$-pole moment. For example, in dipole-conserving systems, the skin effect maximizes the quadrupole moment (see Fig 1(b)).
In their seminal works, Hatano and Nelson studied the effect of disorder on the non-Hermitian skin effect in the non-interacting case Hatano and Nelson (1996); *Hatano1997; *Hatano1998NonHermitian. They found that in open boundary conditions, the system undergoes a transition between two localized phases as a function of disorder strength. At weak disorder, the skin effect persists and wavefunctions are localized at a boundary; at strong disorder, the skin effect is overcome by Anderson localization, and wavefunctions are localized at random positions. When the boundary conditions are periodic, the non-reciprocal hopping creates a persistent current, and the skin effect phase becomes delocalized. In contrast, the Anderson-localized phase remains largely unaffected, since only a few localized modes are impacted by the change in boundary conditions (see Fig. 1(a)). Recently, the stability of this transition has been numerically investigated in interacting systems, which appear to have a non-Hermitian MBL phase Hamazaki et al. (2019); Zhai et al. (2020); Suthar et al. (2022); Ghosh et al. (2022); De Tomasi and Khaymovich (2024); Mák et al. (2024); Roccati et al. (2024).
Figure 1: (a) The charge skin effect (OBC) and persistent charge current (PBC) give way to Anderson localization as disorder strength is increased (b) The dipole skin effect (OBC) and persistent dipole currents (PBC) are stable to disorder.
Here we investigate the impact of disorder on systems that exhibit the non-Hermitian multipole skin effect. First, we generalize the Hatano-Nelson argument to the interacting case and show that, in the presence of U(1) charge conservation, there is a transition between a non-Hermitian skin effect phase and an MBL phase. In the dynamics, this corresponds to a volume-to-area law entanglement phase transition. Next, we consider multipole-conserving systems and demonstrate that, for any amount of disorder, the multipole skin effect remains stable. This also implies that eigenstates are always delocalized for periodic boundary conditions.
Charge Skin Effect.—We begin by reviewing how disorder modifies the non-Hermitian skin effect for U$(1)$ charge Hatano and Nelson (1996); Brouwer et al. (1997); Feinberg and Zee (1999) through the lens of the multipole skin effect Gliozzi et al. (2024). This approach clarifies the role of interactions and generalizes to more exotic skin effects for dipoles and higher multipoles. The minimal model of non-reciprocal charge hopping in a disordered potential is the one-dimensional Hatano-Nelson Hamiltonian Hatano and Nelson (1996),
$H_{\text{HN}}(g)=\sum_{j=-L/2}^{L/2}(e^{g}c^{\dagger}_{j+1}c_{j}+e^{-g}c^{\dagger}_{j}c_{j+1})+\sum_{j}h_{j}n_{j},$ (1)
where $c^{\dagger}_{j}$ is a fermion creation operator at site $j$, $h_{j}\in[-W,W]$ are random fields, $n_{j}=c^{\dagger}_{j}c_{j}$, $g$ sets the strength of the non-Hermiticity, and $L$ is the system size.
In the Hermitian limit $g=0$, all single-particle eigenstates are Anderson localized for any disorder Abrahams et al. (1979). In the clean limit ($W=0,g\neq 0$), the model exhibits the non-Hermitian skin effect: states localize at one edge in open boundary conditions (OBC) and delocalize in periodic boundary conditions (PBC) Hatano and Nelson (1996). Away from these limits, the interplay between disorder and non-reciprocity leads to a transition. For OBC, the transition separates a skin effect phase at weak disorder from Anderson localization at strong disorder, while for PBC it pits localization against delocalization.
Figure 2: Non-interacting Hatano-Nelson model, $W=4$. (a) Phase diagram as a function of disorder strength $W$ and non-Hermiticity $g$. (b) An extensive dipole moment in many-body OBC eigenstates indicates the charge skin effect. (c) A finite fraction of PBC energies with a nonzero imaginary part, $f_{imag}$, indicates delocalization (insets: sample single-particle spectra), (d) The imbalance $\mathcal{I}(t)$, after evolving a Néel state in PBC, does not go to zero in the localized phase at long times. Quantities with an overline are averaged over disorder and mid-spectrum eigenstates.
This transition can be understood by making use the similarity transformation Hatano and Nelson (1996)
$S^{-1}c^{\dagger}_{j}S=e^{-gj}c^{\dagger}_{j},\quad S^{-1}c_{j}S=e^{gj}c_{j},$ (2)
which maps the non-Hermitian model in OBC to its Hermitian version: $S^{-1}H_{\text{HN}}(g)S=H_{\text{HN}}(0)$. The existence of this transformation automatically implies that the OBC spectrum of $H_{\text{HN}}(g)$ is real. Furthermore, the right eigenstates of the non-Hermitian model can be obtained from the corresponding Hermitian ones: |$\ket{\psi_{g}}=S\ket{\psi_{0}}$. Single-particle states are created by the operators
$\eta^{\dagger}_{\alpha}(0)=\sum_{j}\phi_{\alpha}(j)c^{\dagger}_{j},\quad\eta^{\dagger}_{\alpha}(g)=S\eta^{\dagger}_{\alpha}(0)=\sum_{j}e^{gj}\phi_{\alpha}(j)c^{\dagger}_{j},$ (3)
where the orbital $\phi_{\alpha}(j)\sim e^{-|j-j_{\alpha}|/\xi}$ is exponentially localized around a site $j_{\alpha}$ with localization length $\xi$  [^1] In the non-Hermitian case, $\eta^{\dagger}_{\alpha}(g)$ is only a creation operator for right eigenstates, such that $|{\psi_{g}^{(\alpha)}}\rangle=\eta^{\dagger}_{\alpha}(g)|{0}\rangle$, and left eigenstates are created by a different operator, $\langle{\psi_{g}^{(\alpha)}}|=\langle{0}|\zeta_{\alpha}(g)$. Thus, when $g<1/\xi$, the non-Hermitian orbitals remain Anderson localized at an arbitrary distance from the boundary, while when $g>1/\xi$ they are localized on the right edge of the system and exhibit the skin effect. Figure 2(a) shows the phase diagram of this system as a function of $g$ and $W$. The localization length is energy-dependent, creating an intermediate region where localized and skin states are separated by a mobility edge. Moreover, at weak (strong) disorder, $\xi\sim 1/W^{2}$ ($\xi\sim 1/\log(W)$Derrida, B. and Gardner, E. (1984); Anderson (1958), leading to a critical non-reciprocity $g_{c}\sim W^{2}$ ($g_{c}\sim\log(W)$).
To establish an order parameter for the transition, we notice that the transformation in Eq. (2) is generated by
$S=e^{g\sum_{j}jn_{j}}=e^{gP},$ (4)
where $P$ is the total dipole moment operator. At large $g$, $S$ tends to project onto states with the largest dipole moment Gliozzi et al. (2024). For single-particle states, this recovers the standard skin effect: clustering charge on the boundary maximizes the dipole moment. However, this map also naturally applies to many-body eigenstates, where the Pauli exclusion principle prevents multiple fermions from localizing at the boundary Lee et al. (2020); Alsallom et al. (2022). We thus define the U(1) charge skin effect more broadly as the tendency of eigenstates to maximize their dipole moment. In Fig. 2(b), the dipole moment renormalized by its maximum value serves as an order parameter for the OBC transition: in the Anderson phase it vanishes, whereas in the skin effect phase it does not.
Now, we turn to the of case of periodic boundary conditions, in which the skin effect phase gives way to delocalization. Heuristically, if the barrier created by the edge is removed, the similarity transformation $S=e^{gP}$ pumps charge preferentially in one direction, leading to a persistent current $j\sim\partial_{t}P$. Without disorder, the energies of delocalized states form a loop in the complex plane: right-movers have a positive imaginary part and are amplified over time, while left-movers have a negative imaginary part and are suppressed. On the other hand, at strong disorder, eigenstates are Anderson localized and therefore cannot be affected by boundary conditions. As a result, the energy of a localized eigenstate under PBC must be the same as its energy under OBC, and hence real. The delocalization transition is therefore associated with a real-to-complex transition of the PBC energy spectrum. The fraction of eigenvalues with a non-zero imaginary part, shown in Fig. 2(c), is therefore a proxy for the transition in PBC. In the Anderson phase, $f_{\text{imag}}\rightarrow 0$, whereas in the delocalized phase, $f_{\text{imag}}\neq 0$.
Finally, we can also identify the localization transition using the non-unitary dynamics defined by
|$\ket{\psi(t)}=\frac{e^{-iH_{\text{HN}}(g)t}\ket{\psi(0)}}{\lVert e^{-iH_{\text{HN}}(g)t}\ket{\psi(0)}\rVert}.$ (5)
Under PBC, time evolution is frozen in the Anderson localized phase, while a persistent unidirectional current carries charge in the skin effect phase. We contrast these behaviors by starting with a Néel state, |$\ket{\psi(0)}=\prod_{i\in 2\mathbb{Z}}c^{\dagger}_{i}\ket{0}$, and tracking the imbalance of charge over time, ($𝜓𝑡\mathcal{I}(t)=\sum_{i}(-1)^{i}\expectationvalue{n_{i}}{\psi(t)}$. In Fig. 2(d), we show the long time limit of the imbalance, which decays to zero for $g>g_{c}$ since the charges are mobile, but remains finite for $g<g_{c}$.
Interacting case.—Having benchmarked the dipole moment as an order parameter for the transition, we now turn to the interacting case. The interacting Hatano-Nelson model is defined as Zhang et al. (2022); Kawabata et al. (2022); Hamazaki et al. (2019)
$H(g)=H_{\text{HN}}(g)+V\sum_{j}n_{j}n_{j+1}.$ (6)
As with any interaction that is diagonal in the occupation basis, the similarity transformation in Eq. (2) continues to map $H(g)$ to $H(0)$ in OBC.
For sufficiently strong disorder, the interacting Hermitian system is believed to host an MBL phase Oganesyan and Huse (2007); Pal and Huse (2010); Luitz et al. (2015); Bera et al. (2015); Devakul and Singh (2015); Sierant et al. (2020); Vardhan et al. (2017); De Tomasi et al. (2017); Doggen et al. (2018). The eigenstates in the MBL phase are adiabatically connected to those at $V=0$ by a quasi-local unitary transformation, and therefore maintain a similar spatial structure Imbrie (2016); Ros et al. (2015); Huse et al. (2014). However, these states cannot be expressed in terms of single-particle orbitals as in Eq. (3), so it is convenient to work with conserved operators instead. In the non-interacting case, the local integrals of motion (LIOMs) associated with the right OBC eigenstates of Eq. (6) are given by
$\mathcal{O}^{(V=0)}_{\alpha}(0):=$ $\eta^{\dagger}_{\alpha}(0)\eta_{\alpha}(0)=\sum_{i,j}\phi_{\alpha}(i)\phi_{\alpha}^{*}(j)c^{\dagger}_{i}c_{j}$ (7)
$\mathcal{O}^{(V=0)}_{\alpha}(g):=$ $\eta^{\dagger}_{\alpha}(g)\eta_{\alpha}(g)=\sum_{i,j}e^{g(i+j)}\phi_{\alpha}(i)\phi_{\alpha}^{*}(j)c^{\dagger}_{i}c_{j}.$
Since $\phi_{\alpha}(i)\phi_{\alpha}^{*}(j)\sim e^{-\text{max}(|i-j_{\alpha}|,|j-j_{\alpha}|)/\xi}$, the non-Hermitian LIOMs are randomly localized for $g<1/\xi$ and edge localized for $g>1/\xi$, indicating the Anderson and skin effect phases, respectively.
Figure 3: Interacting charge-conserving Hatano-Nelson model, $V=1,W=8$ (a) Eigenstate dipole moment $P$ divided by its maximum value. For $g<g_{c}$ the system is in an MBL phase and $\overline{P}/P_{max}\rightarrow 0$, while this limit is finite in the skin effect phase for $g>g_{c}$. The dipole moment $P$ is averaged over disorder and mid-spectrum eigenstates. (b) Volume-to-area law entanglement transition in long-time steady states of the dynamics (inset: entanglement entropy fluctuations peak at the transition). The bipartite von Neumann entropy $S_{L/2}(t)$ is averaged over disorder and random initial states.
In an MBL phase with $V>0$, LIOMs are locally dressed by the weak interactions:
$\mathcal{O}^{(V)}_{\alpha}(0)$ $=\sum_{i,j}\tilde{\phi}_{\alpha}(i,j)c^{\dagger}_{i}c_{j}$ (8)
$+\sum_{i,j,k,l}\tilde{\phi}_{\alpha}(i,j,k,l)c^{\dagger}_{i}c^{\dagger}_{j}c_{k}c_{l}+\ldots.$
These operators reduce to Eq. (7) as $V\rightarrow 0$ and fall off exponentially away from the localization center $j_{\alpha}$, such that $\tilde{\phi}_{\alpha}(\{x_{i}\})\sim e^{-\text{max}(|x_{i}-j_{\alpha}|)/\xi}$ ($\xi$ is not necessarily the same as in the non-interacting case). All LIOMs commute with the total U$(1)$ charge operator $N=\sum_{j}n_{j}$  [^2] Since both the LIOMs and the charge commute with the Hamiltonian, if $[\mathcal{O}_{\alpha},N]\neq 0$, then there would be a degenerate subspace of states, which is almost impossible due to the onsite disorder. See also Ref. Potter and Vasseur, 2016.. Acting with the similarity transformation, we obtain the corresponding non-Hermitian (right-eigenstate) LIOMs:
$\mathcal{O}^{(V)}_{\alpha}(g)=\sum_{n=1}\sum_{x_{1},\ldots,x_{2n}}\!\!e^{g(x_{1}+\ldots+x_{2n})}\tilde{\phi}_{\alpha}(x_{1},\ldots,x_{2n})c^{\dagger}_{x_{1}}\ldots c_{x_{2n}}.$ (9)
In any given term of this expansion, the coefficient of the operator scales as $e^{g\max(x_{i})}e^{-\text{max}(|x_{i}-j_{\alpha}|)/\xi}$, so localization occurs when $g<1/\xi,$ and the skin effect occurs when $g>1/\xi$. Thus, the many-body similarity transformation also explains the existence of non-Hermtian MBL, which was observed in Ref. Hamazaki et al., 2019. We confirm the transition numerically by using the dipole moment as an order parameter (see Fig. 3(a)). When the system is periodic, the skin effect is replaced by a persistent current, and the transition is between MBL and delocalized eigenstates. This transition had been observed numerically in PBC Hamazaki et al. (2019); Mák et al. (2024), though its connection to the skin effect and its explanation in terms of LIOMs were not noted.
Interestingly, although the transition connects two localized phases in an OBC system, interactions introduce a crucial difference in the dynamics of quantum entanglement  [^3] We quantify entanglement via the half-partition von Neumann entanglement entropy, $S_{L/2}(t)=-\text{Tr}[\rho_{L/2}(t)\log\rho_{L/2}(t)]$, with $\rho_{L/2}(t)$ the bipartite reduced density matrix of the evolved state $|\psi(t)\rangle$, starting from the Néel state.. In the MBL phase, interaction-induced dephasing drives the system toward a volume-law-entangled state at long times. Conversely, in the skin-effect phase, the state is projected onto its component with maximal dipole moment, resulting in area-law entanglement Gliozzi et al. (2024). The MBL-to-skin-effect transition in OBC therefore induces an entanglement transition from volume- to area-law behavior Kawabata et al. (2023), as shown in Fig. 3(b).
Multipole Skin Effect.—We now turn to the skin effect in systems that conserve both a U$(1)$ charge, like the total particle number $N=\sum_{j}n_{j}$, and its dipole moment $P=\sum_{j}jn_{j}$. A dipole-conserving generalization of the Hatano-Nelson model is
$H_{5}(g_{0},g_{1})=\!\!\!\!\sum_{r\in\{0,1\},\,j}\!\!\!\!t_{r}(e^{g_{r}}c^{\dagger}_{j}c_{j+1}c_{j+2+r}c^{\dagger}_{j+3+r}+e^{-g_{r}}\times\mathrm{h.c.}),$ (10)
where the kinetic terms describe imbalanced dipole hopping as depicted in Fig. 1(b). We consider the model at half-filling ($N=L/2$) and in the largest dipole sector ($P=0$), and we initially concentrate on the clean limit. Without disorder, this system exhibits a dipole skin effect, whereby dipoles are preferentially pumped to one boundary Gliozzi et al. (2024). Unlike the conventional (charge) skin effect, which clusters charges at one boundary, the dipole skin effect pushes charges symmetrically outwards to both edges of an open chain. In a system without boundaries, this leads to a persistent dipole current and delocalized states (schematically shown in Fig. 1(b)).
For the special case of Eq. (10) with $g_{1}=3g_{0}/2$, the dipole skin effect can be understood via a similarity transformation analogous to Eq. (4):
$S=e^{g_{0}\sum_{j}j^{2}n_{j}/4}=e^{g_{0}\mathcal{Q}/4},$ (11)
where $\mathcal{Q}$ is the quadrupole moment of the charge distribution. This map transforms $H_{5}(g_{0},3g_{0}/2)\mapsto H_{5}(0,0)$ via $(c^{\dagger}_{j},\,c_{j})\mapsto(e^{-g_{0}j^{2}/4}c^{\dagger}_{j},\,e^{g_{0}j^{2}/4}c_{j})$, enforcing a real spectrum in OBC. Much like the charge skin effect produces eigenstates with an extensive dipole moment, the dipole skin effect leads to eigenstates with an extensive quadrupole moment. Moreover, these properties hold even when Eq. (11) does not map $H_{5}(g_{0},g_{1})$ to a Hermitian Hamiltonian Gliozzi et al. (2024).
Having discussed the skin effect in the clean limit, we return to systems with a disordered potential,
$H_{\text{dip}}(g):=H_{5}(g,3g/2)+\sum_{j}h_{j}n_{j}.$ (12)
At strong enough disorder  [^4] For the chosen disorder strength $W=5.5$, we verify that the dynamics is frozen (see the inset of Fig. 4(c)) and that the level-spacing distribution is Poissonian. See Supplementary Information for additional plots, the model is MBL when $g=0$ Herviou et al. (2021). In analogy with the charge-conserving case, we investigate the existence of a putative transition between MBL and the dipole skin effect at some $g_{c}\neq 0$. We provide evidence that this is not the case: any amount of non-Hermiticity spoils MBL and induces the dipole skin effect. In particular, this implies that the system with PBC is delocalized for any disorder strength.
Figure 4: Dipole-conserving Hatano-Nelson model, $W=5.5$, $t_{1}=0.3t_{0}$. (a) The dipole skin effect manifests as extensive quadrupole moment in OBC eigenstates, and is present for any $g>0$ (b) The fraction of PBC energies with nonzero imaginary part tends to a finite value, indicating delocalization (inset: sample energy spectrum). Both quantities are averaged over disorder and mid-spectrum eigenstates. (c) Time evolution of imbalance starting from the Néel state in PBC with $L=24$. Imbalance remains finite at long times for $g=0$ (inset: different system sizes), while it decays for $g>0$.
Our argument is very similar to the charge-conserving case (c.f. Eqs. (8)–(9)) and relies on applying the similarity transformation to the Hermitian, dipole-conserving LIOMs. In the $g=0$ MBL phase, each state can be labeled by an extensive number of LIOMs,
$\mathcal{O}^{\text{dip}}_{\alpha}(0)=\sum_{i+k=j+l}{\varphi}_{\alpha}(i,j,k,l)c^{\dagger}_{i}c_{j}c_{k}c^{\dagger}_{l}+\ldots,$ (13)
where $\varphi_{\alpha}(\{x_{i}\})\sim e^{-\text{max}(|x_{i}-x_{\alpha}|)/\xi}$ for some localization center $x_{\alpha}$, the dots indicate higher-order terms, and dipole symmetry is explicitly enforced in each term  [^5] Diagonal terms like $\sum_{i}\varphi_{\alpha}(i)n_{i}$ are also included in the sum, as they conserve dipole moment. Applying the map, we obtain the (right-eigenstate) LIOMs of the non-Hermitian system  [^6] Right eigenstates define the non-Hermitian LIOMs, so creation and annihilation operators are scaled by the same factor, as in Eq. (7).:
$\mathcal{O}^{\text{dip}}_{\alpha}(g)=\sum_{i+k=j+l}e^{g(i^{2}+j^{2}+k^{2}+l^{2})/4}{\varphi}_{\alpha}(i,j,k,l)c^{\dagger}_{i}c_{j}c_{k}c^{\dagger}_{l}+\ldots.$ (14)
Although $\varphi_{\alpha}$ is exponentially localized around a random lattice site $x_{\alpha}$, the prefactor $e^{g(i^{2}+j^{2}+k^{2}+l^{2})/4}$ is super-exponentially localized at the outer edges of the chain. This latter factor always wins out, and for any $g>0$, the non-Hermitian LIOMs are pushed symmetrically to the boundaries of an OBC system. Consequently, the dipole skin effect always prevails over MBL.
Heuristically, there is a competition between two tendencies: localizing charges randomly and localizing dipoles on the boundary. However, charges are zero-dimensional objects while dipoles are one-dimensional. The dipole skin effect can thus arise both from fixed-length dipoles moving to the edge and from the boundary dipoles themselves growing in length. The latter process is unique to dipoles, and provides extra pathways to maximize the quadrupole moment. It is this additional localization mechanism that allows the dipole skin effect to dominate over the localization of charge.
We confirm this numerically by plotting the average quadrupole moment of the eigenstates of $H_{\text{dip}}(g)$ in Fig. 4(a): the system develops an extensive quadrupole moment that grows with system size for any $g\neq 0$. In PBC, the skin effect becomes a persistent dipole current that delocalizes the system. To see this, we plot the fraction of complex energies $f_{\text{imag}}$ in Fig. 4(b). In agreement with our analytical argument, the density of complex energies tends to a finite value with increasing system size. We also investigate the out-of-equilibrium dynamics through the imbalance $\mathcal{I}(t)$ starting from the previously defined Néel state, shown in Fig. 4(c). After a transient, the dynamics is frozen for $g=0,$ but mobile for any amount of non-Hermiticity.
The primacy of the dipole skin effect over MBL also extends to the broader family of multipole skin effects Gliozzi et al. (2024). By analogy to the charge and dipole skin effects, the $m$-pole skin effect involves $m$-poles preferentially hopping to one side to generate an $(m+1)$-pole moment. The analogous similarity transformation is generated by $S=e^{f(g)\sum_{j}j^{m+1}n_{j}}$, and Hermitian LIOMs, which can at most be exponentially localized Imbrie (2016); Thomson and Schiró (2018); Chandran et al. (2015); Ros et al. (2015); Huse et al. (2014); Serbyn et al. (2013); Rademaker et al. (2017), are pushed away from their localization centers by the non-Hermiticity. If $m\geq 1$, the multipole skin effect overcomes disorder in OBC, and delocalization ensues in PBC.
Discussion.—We have seen that the fate of localization in disordered, non-Hermitian systems strongly depends on the underlying symmetries. In non-reciprocal systems with a conserved U$(1)$ charge, there is a transition between localization and the non-Hermitian skin effect. Under PBC, this becomes a delocalization transition, as the skin effect gives way to a persistent current. However, when a multipole moment of the charge is also conserved, the skin effect always dominates over localization. This implies that a PBC system with multipole conservation, which imposes kinetic constraints that hinder particle motion Sala et al. (2020); Feldmeier et al. (2020), and strong disorder, which also tends to freeze particles, is nevertheless delocalized as long as some degree of non-reciprocity is present.
One consequence of our results is that multipole skin effects can be stabilized even when multipole conservation is explicitly broken by charge hopping: there exists a regime where disorder localizes the charge while multipoles remain delocalized. Relatedly, it would be interesting to investigate possible transitions between skin effects for different multipoles, like dipoles and charges. The effects of discrete and non-Abelian symmetries on the disordered skin effect should also be explored. Finally, moving beyond one-dimensional systems remains an important avenue for further study.
Acknowledgements.—J.G. and T.L.H. acknowledge support from the US Office of Naval Research MURI grant N00014-20-1-2325. G.D.T. acknowledges support from the EPiQS Program of the Gordon and Betty Moore Foundation and the hospitality of MPIPKS Dresden, where part of the work was done.
Supplementary Material
SM 1: Skin effects without the similarity transformation to a Hermitian Hamiltonian
In the main text, we focused on non-Hermitian systems with a real spectrum in open boundary conditions. In these cases, a similarity transformation maps the non-Hermitian Hamiltonian to a Hermitian one. For both the charge skin effect and the dipole skin effect, our arguments about stability to disorder were based on applying this similarity transformation to obtain non-Hermitian eigenstates from their Hermitian counterparts.
Nevertheless, it was argued in Ref. Gliozzi et al., 2024 that even Hamiltonians with a complex OBC spectrum exhibit the same skin effect phenomenology. Here, we show that this is also true for these Hamiltonians with onsite disorder. Concretely, we consider non-reciprocal models that do not admit a similarity transformation to a Hermitian system, and numerically show that they have the same key features and localization transitions as the models in the main text.
For U(1) charge conservation, we can “break” the similarity transformation by adding a next-nearest neighbor hopping term to the Hatano-Nelson model of Eq. (1):
$\tilde{H}_{\text{charge}}=-\sum_{j}[e^{g}(c^{\dagger}_{j+1}c_{j}+c^{\dagger}_{j+2}c_{j})+e^{-g}\times\mathrm{h.c.}]+\sum_{j}h_{j}n_{j},$ (S1)
where again the disordered potential $h_{j}\in[-W,W]$. Even in open boundary conditions, this Hamiltonian does not have a real spectrum for $g>0$, and therefore does not admit a similarity transformation to a Hermitian model. Nevertheless, we numerically verify in Fig. S1(a) that it still has a transition between an Anderson localized phase and a phase exhibiting the charge skin effect. In Fig. S1(d), show an approximate data collapse when the non-reciprocity is rescaled by $L^{1/4}$.
In the dipole-conserving case, we can “break” the similarity transformation by choosing different imbalances $g_{0}$ and $g_{1}$ for the range-four and range-five dipole hopping terms in Eq. (12):
$\tilde{H}_{\text{dip}}(g)=H_{5}(g,2g)+\sum_{j}h_{j}n_{j},$ (S2)
This Hamiltonian has a complex spectrum in open boundary conditions and does not admit a similarity transformation to a Hermitian model. Its eigenstates cannot therefore be obtained by applying $S=e^{g\mathcal{Q}/4}$ to a set of Hermitian eigenstates. However, we numerically verify in Fig. S1(b) that it still exhibits the dipole skin effect for any $g>0$, as in the case where the map works. In Fig. S1(d), we also notice an approximate data collapse when the non-reciprocity is rescaled by $\sqrt{L}$.
In general, our predictions for localization transitions or lack thereof should also be valid for systems in which the similarity transformation does not work, as long the direction of non-reciprocity remains the same. Fundamentally, all multipole skin effects are driven by an imbalanced hopping of $m$-poles. Independently of whether a similarity transformation to a Hermitian Hamiltonian exists, such hopping creates a current of $m$-poles that tends to maximize the $(m+1)$-pole moment in an open system.
Figure S1: Skin effect order parameters for models without a similarity transformation (a) The charge-conserving Hatano-Nelson model with next-nearest neighbor hoppings, $\tilde{H}_{\text{charge}}(g)$, has a transition between Anderson localization and the skin effect as measured by eigenstate dipole moment ($W=4$). (b) The dipole-conserving model $\tilde{H}_{\text{dip}}(g)$ appears to exhibit the dipole skin effect for any $g>0$, as measured by eigenstate quadrupole moment. The data collapse reasonably well when non-reciprocity is rescaled by (c) $L^{1/4}$ ($g_{c}=0.38$) for the charge-conserving case and (d) $\sqrt{L}$ ($g_{c}=0$) for the dipole-conserving case.
SM 2: Additional evidence for many-body localization in the Hermitian dipole-conserving model
In the main text, we considered a dipole-conserving model with strong enough onsite disorder $W=5.5$ such that the Hermitian version, $H_{\text{dip}}(g=0)$ is many-body localized. We showed dynamical evidence for MBL in Fig. 4(c), where the long-time imbalance starting from the Néel state is seen to saturate at a nonzero value when $g=0$. Here we present further spectral evidence that the Hermitian system with $W=5.5$ is indeed many-body localized, focusing on the symmetry sector containing the Néel state.
The first signature we consider is the energy level statistics. Chaotic Hamiltonians exhibit level repulsion, and midspectrum energies are distributed according to the Gaussian Orthogonal Ensemble (GOE). On the other hand, MBL systems do not have revel repulsion, and midspectrum energies have a Poissonian distribution. The proximity to either limit can be quantified by ratio of consecutive energy gaps,
$r_{n}=\frac{\min(s_{n},s_{n+1})}{\max(s_{n},s_{n+1})},\quad s_{n}=E_{n+1}-E_{n}.$ (S3)
The average value of $r_{n}$ (over disorder and eigenstates) is called the r-statistic $\overline{r}$. In MBL systems, we expect a value close to $r_{\text{Poisson}}\approx 0.386$, while in a chaotic system we expect $r_{\text{GOE}}\approx 0.531$. In Fig. S2(a), the r-statistic at $W=5.5$ is very close to Poissonian, indicating an MBL phase.
The second signature we consider is the entanglement entropy of mid-spectrum eigenstates. Concretely, we consider the von Neumann entanglement entropy for a bipartition of our system, $S_{L/2}=-\text{Tr}[\rho_{L/2}\log\rho_{L/2}]$, where $\rho_{L/2}$ is the reduced density matrix of an eigenstate on half of the chain. In a chaotic system, these states are expected to be thermal and therefore have a volume-law entanglement, $S_{L/2}\sim L$. In an MBL system, midspectrum eigenstates are non-thermal and area-law entangled, $S_{L/2}\sim\mathcal{O}(1)$. In Fig. S2(b), the entanglement entropy at $W=5.5$ appears to not scale with system size, indicating area-law entanglement.
Figure S2: Spectral evidence of many-body localization in the Hermitian, dipole conserving model $H_{\text{dip}}(0)$. (a) The spectral statistics of midspectrum energy levels are Poissonian at $W=5.5$, (b) The half-chain entanglement entropy of midspectrum eigenstates is area law scaling at $W=5.5$.
SM 3: Localization to skin effect transition in a model with hidden U$(1)$ symmetry
We have argued that non-reciprocal U(1) charge hopping permits a transition between disorder-induced localization and the skin effect. On the other hand, non-reciprocal dipole hopping, which conserves both U(1) charge and its dipole moment, always leads to the dipole skin effect and never many-body localization.
Here we discuss a special case that naively appears to forbid a localization transition but actually hosts one because of a hidden U(1) symmetry. Consider a dipole-conserving Hamiltonian with only the minimal-range hopping term:
$H_{4}(g)=\sum_{j}(e^{g}c^{\dagger}_{j}c_{j+1}c_{j+2}c^{\dagger}_{j+3}+e^{-g}\times\mathrm{h.c.})+\sum_{j}h_{j}n_{j}.$ (S4)
Let us restrict our attention to the symmetry sector at half-filling $N=L/2$ and dipole moment $P=0$. Inside this sector, the limited range of the dipole hopping imposes a strong kinetic constraint: many states are not connected to each other by the Hamiltonian. In fact, the sector itself fragments into exponentially many disconnected Krylov subsectors, and the largest of these is exponentially small in system size ($D_{\text{largest}}/D_{\text{total}}\sim e^{-L}$ala et al. (2020); Moudgalya et al. (2020). While in the main text we avoided this (strong) Hilbert space fragmentation by adding a longer-ranged dipole hopping term, it is also interesting to consider how the skin effect and disorder are manifest in individual Krylov subsectors.
The largest Krylov subsector is integrable and can be mapped to a free-fermion chain of length $L/2$ Moudgalya et al. (2020). The mapping involves grouping sites $12k-$ 1 and $2k$ of the original lattice into a new composite site. After this rewriting, the integrable subsector is the one where each pair of sites contains only a single occupied state. In other words, only the computational basis states |:=|$\ket{\rightarrow}:=\ket{01}$ and |:=|$\ket{\leftarrow}:=\ket{10}$ are allowed on every (odd-even) pair of sites. These two states represent a local dipole and a local anti-dipole, respectively. For a composite site $k:=(2k-1,2k)$, we can define dipole creation and annihilation operators, $d^{\dagger}_{k}:=c_{2k-1}c^{\dagger}_{2k}$ and $d_{k}:=c^{\dagger}_{2k-1}c_{2k}$. These operators transform a local dipole into a local anti-dipole and vice versa. When acting on the integrable subspace, they obey fermionic anticommutation relations, so they create genuine local quasiparticle excitations. Furthemore, if we define a number operator for these local dipoles, $n^{d}_{k}=d^{\dagger}_{k}d_{k}$, then the total dipole operator in the subsector can be written as
$P=\sum_{j=1}^{L}jn_{j}=\sum_{k=1}^{L/2}n^{d}_{k},$ (S5)
where the first expression is on the original lattice and the second expression uses the composite degrees of freedom. In this sector, the conservation of dipole moment acts as an ordinary U(1) symmetry with local charges.
Since only the odd terms of (S4) act nontrivially on the subspace, we can project the Hamiltonian into the subsector using the composite degrees of freedom:
$H^{\text{sub}}_{4}(g)=\sum_{k}(e^{g}d^{\dagger}_{k+1}d_{k}+e^{-g}d^{\dagger}_{k}d_{k+1})+\sum_{k}\tilde{h}_{k}n^{d}_{k},$ (S6)
where $\tilde{h}_{k}:=h_{2k}-h_{2k-1}$ is an onsite potential derived from the action of the original disorder on the new degrees of freedom, and we have ignored an overall constant. This Hamiltonian is identical to the Hatano-Nelson model, though with the interpretation that the charges are actually local dipoles. Moreover, $H^{\text{sub}}_{4}(g)$ maps to its Hermitian counterpart under the similarity transformation generated by $S=e^{g\sum_{k}kn^{d}_{k}}$, which maximizes the dipole moment of the local dipoles. As a result, it hosts a transition between Anderson localization and the “charge skin effect”, where both types of localization involve local dipole quasiparticles.
Such a transition can only occur if the dipoles are local quasiparticles as in this Krylov subsector. Generally, dipoles are 1D objects and their localization (via the dipole skin effect) dominates over the localization of 0D charge in a possible Anderson/MBL phase. However, if dipoles can be mapped to local charges, then, like charged particles, their Anderson localization competes with the skin effect. Nevertheless, this is a highly fine-tuned situation. For the dipole moment to decompose into a sum of local densities as in Eq. (S5), the Hilbert space must be truncated to forbid states containing longer dipoles, and the Hamiltonian must be carefully tuned to preserve these states. For example, even $H_{4}(g)$ does not exhibit this behavior in general, and only a vanishingly small subsector of the Hilbert space hosts a transition.
The hidden U(1) symmetry manifests in another way: the dipole skin effect is more weakly localized than usual inside the Krylov subsector. The dipole skin effect maximizes the quadrupole moment, which generically scales as $L^{3}$ for many-body states, leading to the super-exponential localization of LIOMs discussed in the main text. In this restricted sector, on the other hand, any quadrupole must be formed by local dipoles, and the largest quadrupole moment, associated with the state |=|$\ket{\leftarrow\leftarrow\ldots\rightarrow\rightarrow}=\ket{1010\ldots 0101}$, only scales as $L^{2}$. This weaker scaling implies only exponential localization of LIOMs from the skin effect, and thus allows a localization transition.